OpenStudy (anonymous):

The value of one trigonometric function is give 0

OpenStudy (amistre64):

cos = x/r sec = r/x sin = y/r csc = r/y tan = y/x cot = x/y

OpenStudy (amistre64):

you have x=4167 and r = 1000 y = sqrt(1000^2 - 4167^2)

OpenStudy (amistre64):

sorry, r = 10000 y = sqrt(10000^2 - 4167^2)